OpenStudy (anonymous):

can someone order these from greatest to least please 0.25, 2/5, 32%, 7/14 If anyone could help that would be great!

OpenStudy (bloomlocke367):

try putting them all in percents

OpenStudy (bloomlocke367):

Do you know how to do that?

OpenStudy (anonymous):

I used to but I forgot

OpenStudy (cwrw238):

multiply by 100 will give you percent

OpenStudy (anonymous):

7/14=0.5 2/5=0.4 32%=32/100=.32 0.25

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

25/100 = 1/4= 35/140 2/5 = 56/140 7/14 = 70/140 32% = 32/100 = 44.8/140 .25, 32%, 2/5 , 7/14 ( from smallest to largest)